[Hallc_running] [New Logentry] RC daily update

hszumila at jlab.org hszumila at jlab.org
Thu Aug 6 15:00:03 EDT 2020

Logentry Text:
Hall C is currently in beam permit, ready for beam after the A/B pass change. The program for tonight is to establish 5 pass beam and push the current to Hall C. 

Summary since yesterday:
* Re-established beam around mid-owl shift. Max reliable current ~10 uA. 
* Accomplished Simona's SoLID Cerenkov test at 1,1.5,2 uA. https://logbooks.jlab.org/entry/3821327
* Tried to run on 3He but could not due to target motion FSD https://logbooks.jlab.org/entry/3821330, https://logbooks.jlab.org/entry/3821413
* Ran the rest of shift and half of day shift on reference cell, all at ~10 uA 
* HMS dipole tripped due to CHL https://logbooks.jlab.org/entry/3821560. We continued running on SHMS during HMS recovery. 
* Ethan fixed the 3He FSD problem by reverting the target ladder position. https://logbooks.jlab.org/entry/3821589 Target is now 1mm lower.

Summary is here: https://hallcweb.jlab.org/wiki/index.php/RC_Update_and_Run_Plan_for_Aug_6,_2020#General_updates

We will have an RC meeting today at 4:15 pm. This is how to connect via bluejeans: https://bluejeans.com/861439920


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3821613
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200806/3e8db6a7/attachment.html>

More information about the Hallc_running mailing list